OpenStudy (anonymous):

Find the derivative of f(x) = 4/x at x = 2 -4 -1 1 4

OpenStudy (primeralph):



OpenStudy (primeralph):

plug in 2

OpenStudy (anonymous):

so its -1???

OpenStudy (anonymous):

@primeralph

OpenStudy (tkhunny):

This is an application of the power rule. \(f(x) = \dfrac{4}{x} = 4x^{-1}\) Thus \(f'(x) = 4(-1)x^{-2} = -\dfrac{4}{x^{2}}\) Evaluating, then, as \(x = 2\), gives \(f'(2) = -\dfrac{4}{(2)^{2}} = -1\)

OpenStudy (primeralph):

yeah, you're right.

OpenStudy (primeralph):

medal?